Suppose you deposit $12,000 in a bank account bearing 7.3% interest compounded monthly.
How many months would it take (to the nearest month) for your account balance to top $15,000? Use the trial and error method with a calculator to find your answer.


Answer by stanbon(75887)

Suppose you deposit $12,000 in a bank account bearing 7.3% interest compounded monthly.
How many months would it take (to the nearest month) for your account balance to top $15,000? Use the trial and error method with a calculator to find your answer.
-------
A(t) = P(1+(r/n))^(nt)
-----
15000 = 12000(1+(0.073/12))^(12t)
-----
5/4 = 1.00608^(12t)
----
12t*log(1.00608) = log1.25
---
Ans:: 12t = 36.81 months
----
Cheers,
Stan H.
------------